# Conformalized Bayesian Inference, with Applications to Random Partition Models

### v1

@@ 3 Methodology: Conformalized Bayesian Inference
Let $\theta,\theta^{\prime}\in\Theta$. The $\mathcal{D}$-KDE score $s(\theta;\bm{\theta}^{1})$ with hyperparameters $\beta,\gamma>0$ is given by
$s(\theta;\bm{\theta}^{1}):=\frac{1}{S}\sum_{t=1}^{S}\mathcal{K}(\theta,\theta_{t}),\qquad\mathcal{K}(\theta,\theta^{\prime}):=\exp\{-\gamma\mathcal{D}(\theta,\theta^{\prime})^{\beta}\}.$
Intuitively, $s(\cdot;\bm{\theta}^{1})$ acts as a kernel density estimator (KDE), but with the Euclidean norm replaced by the metric $\mathcal{D}$ to handle the possibly complex parameters in $\Theta$ (Schölkopf and Smola,, 2001; Shawe-Taylor and Cristianini,, 2004). Under certain conditions on $(\Theta,\mathcal{D})$, one can have $\mathcal{D}(\theta,\theta^{\prime})=\|\phi(\theta)-\phi(\theta^{\prime})\|$ for some isometric embedding $\phi:\Theta\to\mathbb{R}^{E}$ and $E\in\mathbb{N}$, at least for certain $\theta,\theta^{\prime}$ (Schoenberg,, 1935). Thus, the proposed score can be viewed—formally in some cases, loosely in general—as a two-step procedure: first embedding parameters into a latent Euclidean space that preserves the geometry induced by $\mathcal{D}$, and then performing KDE in that space.
While other scoring rules can be devised, $\mathcal{D}$-KDE provides a simple and robust baseline. It relies on a clearly defined metric that encodes a meaningful notion of distance between parameters (a computationally or conceptually valuable feature, depending on the context) and reduces the challenge of defining a density-like score on a complex space $\Theta$ to a straightforward, efficiently implementable KDE-like computation. In particular, evaluating $\mathcal{K}(\theta,\theta_{t})$ is easily parallelized across samples in the calibration and/or training sets, and computation can be further sped up by computing the score, for each calibration sample, using only a random subset of the training samples.
3.1 Point estimation
@@
$\theta_{\star}\in\arg\max_{\theta\in\bm{\theta}^{2}}s(\theta;\bm{\theta}^{1}).$
Because $s(\cdot;\bm{\theta}^{1})$ is constructed to act as a posterior KDE evaluation, $\theta_{\star}$ can be interpreted as a kind of maximum-a-posteriori (MAP) estimator based on this posterior KDE. Of course, this interpretation should be taken with caution, since the posterior may not possess a proper continuous density (as in random partition models, where the posterior over $\Theta$ is a probability mass function and the true MAP corresponds to the partition with the largest mass), although it provides a simple and efficient strategy to summarize the posterior with a single parameter value.
Another useful perspective is to view $\theta_{\star}$ as a posterior expected loss minimizer restricted to the calibration set, with the only difference that the loss $\mathcal{D}$ is transformed via $d\mapsto\exp\{-d\}$ to produce a density-like score. In contexts such as random partition models, where efficient algorithms exist for exhaustive searches over $\Theta$ (Wade and Ghahramani,, 2018; Dahl et al.,, 2022), ideas from that literature can be used to extend the search for the score maximizer beyond the calibration set. Nonetheless, we believe that restricting the maximization to the calibration set remains appealing for several reasons: (i) the scores must in any case be computed (and can be parallelized) for the uncertainty quantification step based on conformal prediction (described later), making this approach more convenient than (possibly sequential) exhaustive search algorithms, when the primary goal is uncertainty quantification and point estimation serves only as a simple preliminary summary of the posterior; (ii) the calibration samples are drawn from the posterior and are therefore likely to achieve high scores in the first place, making exhaustive search potentially redundant; and (iii) the search can be flexibly customized—for example, in the context of clustering, one may prefer focusing on partitions with at most three scientifically interpretable clusters, in which case the score maximization can be restricted to the subset of calibration samples satisfying the desired properties. This allows the procedure to combine data-driven evidence with user-specified structural assumptions in a flexible and interpretable way.
@@
The data are then analyzed by fitting a Gaussian mixture model with a Pitman-Yor process prior on the mixing measure (Pitman and Yor,, 1997; Perman et al.,, 1992; De Blasi et al.,, 2013), using a concentration parameter of 0.03 and a discount parameter of 0.01. A standard marginal Gibbs sampler is applied via the Python package pyrichlet (Selva et al.,, 2025), and after burn-in and thinning, 5,000 training partitions and 1,000 calibration partitions are retained.
@@
Figure 1(b) shows the posterior distribution of $K$, the number of distinct clusters assigned to the 100 data points. The histogram suggests that values 2, 3, and 4 all receive substantial posterior mass, indicating the presence of potential uncertainty in the clustering. Figure 1(c) displays the partition corresponding to the $\mathcal{D}_{VI}$-KDE point estimate (with hyperparameters $\gamma=0.5,\beta=1$). This estimate essentially coincides with the true clustering once the two left-most clusters are merged. Notably, this occurs even though partitions with 3 clusters receive more posterior mass than those with 2 clusters (see Figure 1(b)). This behavior may be due to many sampled partitions with 3 clusters being very close to the point estimator (e.g., differing from it in only a few cluster assignments), or because the VI metric inherently favors parsimonious clusterings by weighting larger clusters more heavily; see Equation (2). A clearer picture will emerge with our upcoming discussion of posterior uncertainty quantification beyond point estimation.
3.2 Credible region
@@
While point estimation provides a useful first summary of the posterior distribution, it necessarily omits the uncertainty encoded in the full posterior. To address this, a key component of our CBI methodology is the construction of a credible posterior subset of $\Theta$ based on the principles of conformal prediction (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Lei et al.,, 2018; Angelopoulos and Bates,, 2023). Very briefly, conformal prediction in its basic form allows one to construct a set-valued prediction for the next observation given the first $N$ samples from an exchangeable sequence, with a precisely defined, finite-sample, and distribution-free coverage guarantee. In our context, the crucial step is to interpret Monte Carlo samples as draws from an exchangeable sequence, and to recast the problem of constructing a posterior credible region as that of predicting a region where a new independent posterior sample would fall with high probability. [^4] We note that, although with entirely different goals (namely, calibrating predictions derived from Bayesian predictive distributions), conformal prediction has already been applied in the Bayesian setting, e.g., by Fong and Holmes, (2021). We now formalize this intuition.
@@
Recall our assumption that the available samples from $\Pi$, and in particular the calibration samples $\bm{\theta}^{2}=(\theta_{S+1},\ldots,\theta_{T})$ (for convenience, let N:=TSN:=T-S), are iid. See Proposition 2 for a theoretical justification of this assumption when the original sampler features Markov dependence, as in MCMC. Define
$\hat{p}(\theta;\bm{\theta}^{2}):=\frac{|\{\theta^{\prime}\in\bm{\theta}^{2}:s(\theta;\bm{\theta}^{1})\geq s(\theta^{\prime};\bm{\theta}^{1})\}|+1}{N+1}.$ (2)
@@
$\mathcal{C}_{1-\alpha}(\bm{\theta}^{2}):=\{\theta\in\Theta:\hat{p}(\theta;\bm{\theta}^{1})\geq\alpha\}.$ (3)
In words, we declare $\theta\in\Theta$ to belong to the credible region $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$ if its $\mathcal{D}$-KDE score is not too low compared to those of the calibration samples—that is, if it is not too “unlikely” under the posterior relative to samples known to be drawn from it. Note that we emphasize the dependence of the credible region on the calibration set but not on the training set. This is because, from now on, the latter is treated as fixed (or, equivalently, we condition on its realization), since the coverage guarantees to be discussed do not depend on the particular scoring rule used to score the calibration samples. The following proposition states precisely the coverage property ensured by $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$.
@@
The set $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$ as defined in Equation (3) satisfies
@@
$\mathbb{P}_{(\bm{\theta}^{2},\theta)\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\theta\in\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\right]\,\equiv\,\mathbb{E}_{\bm{\theta}^{2}\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\Pi\left(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\right)\right]\,\geq\,1-\alpha.$ (4)
Furthermore, for any $\delta\in(0,1)$, its posterior mass $\Pi\left(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\right)$ satisfies
@@
$\left|\Pi\left(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\right)-(1-\alpha)\right|\leq\frac{\max(\alpha,1-\alpha)}{N}+\left|\frac{k_{\alpha,N}}{N}-\hat{\Pi}_{s,N}\big(s_{(k_{\alpha,N})}^{-}\big)\right|+\sqrt{\frac{1}{2N}\ln\left(\frac{2}{\delta}\right)}$ (5)
@@
with $\mathbb{P}_{\bm{\theta}^{2}\overset{\textnormal{iid}}{\sim}\Pi}$-probability at least $1-\delta$, where $k_{\alpha,N}:=\lceil\alpha(N+1)-1\rceil$, $s_{(k)}$ is the $k$-th order statistic of the calibration scores, and $\hat{\Pi}_{s,N}$ is their empirical CDF. Finally, $\hat{p}(\theta;\bm{\theta}^{2})$ as defined in Equation (2) satisfies
@@
$\forall\lambda\in[0,1],\quad\mathbb{P}_{(\bm{\theta}^{2},\theta)\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\hat{p}(\theta;\bm{\theta}^{2})\leq\lambda\right]\leq\lambda.$ (6)
@@
Proposition 1 clarifies the rationale behind defining $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$ as a credible region under the posterior $\Pi$ with coverage $(1-\alpha)\times 100\%$. In particular, Equation (4) ensures that the entire procedure, i.e., scoring the calibration samples and checking whether an independent draw from $\Pi$ falls within $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$, succeeds with probability at least $1-\alpha$. Unsurprisingly, given its connection to conformal prediction, this construction has a frequentist flavor: it yields a distribution-free, finite-sample coverage guarantee under $\Pi$ based on samples from it. From a purely subjective Bayesian standpoint, however, this raises no philosophical tension, since the task is to infer properties of the posterior distribution itself, which exists and is fixed, independently of whether the Bayesian model giving rise to it is meant to learn any true, fixed data-generating process. On the other hand, a key advantage of the conformal prediction approach is that the resulting coverage guarantee is not tied to any property of the posterior distribution. In particular, it does not rely on any specific topological or algebraic structure of $\Theta$, which in many cases of interest may be non-standard, and remains valid regardless of the dimensionality of the posterior support, which in many of our motivating applications is extremely large or even infinite.
@@
As Equation (4) shows, the coverage guarantee can also be interpreted as an expected-mass statement: the posterior mass of $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$ exceeds $1-\alpha$ in expectation, where the randomness arises from the calibration set. Thus, $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})$ is a random subset of $\Theta$ whose expected posterior mass is at least $1-\alpha$. Equation (5) further characterizes the behavior of this random set, showing that, with high probability, its posterior mass concentrates around $1-\alpha$ at rate $N^{-1/2}$ in the calibration size $N$. Note that, unless the posterior distribution of the scores is atomless, the second term on the right-hand side of Equation (5) need not vanish as $N\to\infty$ (while it equals $1/N$ when the score distribution is continuous). This reflects the well-known fact that conformal quantile-based sets achieve coverage at least as high as the nominal level $1-\alpha$, but may not approximate it tightly from above if atoms at specific quantiles exist in the target score distribution.
@@
This notion of coverage differs from the traditional Bayesian definition, which concerns a fixed, non-random credible region $\mathcal{C}$ satisfying $\Pi(\mathcal{C})\geq 1-\alpha$. Such a region, if computable and sufficiently small, would of course be preferable, as it would provide deterministic coverage. However, the very complexity of $\Theta$ that motivates our framework typically precludes analytical construction of such sets. Moreover, even conventional credible regions derived from Monte Carlo samples are inherently random, although this is rarely acknowledged explicitly: for instance, even a simple credible interval for a one-dimensional parameter (e.g., based on empirical upper- and lower-quantiles of Monte Carlo samples from the posterior distribution of that parameter) inherits randomness from the finite Monte Carlo sample size. Or, focusing on our application to random partition models, Wade and Ghahramani, (2018) define a $1-\alpha$ credible region under the posterior on partitions as the smallest $\mathcal{D}_{VI}$ ball centered at the minimum expected $\mathcal{D}_{VI}$ estimator attaining the target posterior mass. In practice, on top of being computationally expensive compared to our parallelizable scoring procedure, this region is computed from MCMC draws, and thus inherits randomness (both in its center and bounds) from the sampling process, though this uncertainty is typically ignored. From this point of view, a key advantage of our conformal construction is that it delivers rigorously defined, finite-sample posterior coverage guarantees, both in expectation (Equation (4)) and with high probability (Equation (5)), that hold for any calibration size $N$, without any assumption on the target posterior distribution.
@@
Moreover, Equation (6) shows that $\hat{p}(\theta;\bm{\theta}^{2})$ can be interpreted as a $p$-value for testing the null hypothesis that $\bm{\theta}^{2}$ and $\theta$ are drawn iid from $\Pi$, since its null distribution is super-uniform. This enables formal hypothesis testing under the posterior, though of a rather different nature from traditional Bayesian procedures such as tests relying on Bayes factors (Jeffreys,, 1998; Kass and Raftery,, 1995). The proposed conformal test has a distinctly frequentist flavor, as it treats the posterior $\Pi$ as an unknown distribution estimated from iid Monte Carlo samples, and the conformal $p$-value provides a classical tool to test a nonparametric null hypothesis on the joint distribution of the calibration samples and a hypothetical new draw from $\Pi$. More generally, if one prefers not to attach a strict testing interpretation to $\hat{p}(\theta;\bm{\theta}^{2})$, this quantity can still be viewed as a normalized measure of the typicality of $\theta$ with respect to $\Pi$: larger values correspond to parameters that rank higher, under the $\mathcal{D}$-KDE score, than most calibration samples. Even under this heuristic interpretation, the conformal analysis based on $\hat{p}(\theta;\bm{\theta}^{2})$ offers a more nuanced characterization of the uncertainty encoded in $\Pi$ than a simple point estimate alone could provide.
@@
An additional appealing feature of the procedure is that it may be applied conditionally on any subregion of $\Theta$ with positive posterior mass. This is particularly useful when one wishes to perform inference while fixing certain aspects of the complex parameter. For example, as already mentioned, in clustering applications one may wish to assess posterior uncertainty conditional on partitions with at most three interpretable clusters. In this case, a valid credible region can be constructed from the calibration scores associated with partitions having at most clusters, with coverage guaranteed relative to the posterior conditioned on the nonzero posterior probability event of three clusters.
@@
On a final note on Proposition 1, it is well known (Angelopoulos and Bates,, 2023) that the coverage guarantee in Equation (4) is invariant to the choice of scoring rule, as it relies solely on the exchangeability between calibration and new samples. Nevertheless, the practical utility of the resulting credible region depends critically on this choice. Arbitrary scores may yield overly large regions (since even the whole $\Theta$ itself trivially satisfies the guarantee) or exclude high concentration areas if the score misaligns with $\Pi$. Designing an informative score is therefore essential. Our distance-based, KDE-like construction offers a principled solution: it exploits (i) the geometry of $\Theta$ induced by $\mathcal{D}$, which helps capture the structure of the posterior’s effective support, and (ii) the inherent concentration of training samples in high-probability regions. In our illustrations, we verify empirically that this choice yields credible regions of reasonable size and consistent posterior alignment in practice.
@@
The concentration inequality in Equation (5) is of independent interest, and its original proof, based on the Dvoretzky–Kiefer–Wolfowitz inequality (Dvoretzky et al.,, 1956; Massart,, 1990), can be found in the Supplementary Material. It is in the spirit of existing results (Vovk,, 2012; Sarkar and Kuchibhotla,, 2023; Lei et al.,, 2018) that study the coverage, under a target distribution $\Pi$, of the conformal set as a random variable conditional on the calibration scores, and establish high-probability bounds for the event of that coverage exceeding $1-\alpha$. Our inequality goes a step further (under the assumption of i.i.d. scores), as it specifies when and at what rate this coverage converges exactly to $1-\alpha$, rather than merely exceeding it. As already noted, the bound includes a term that may not vanish asymptotically in the presence of atoms in the score distribution; when $\Pi$ is continuous, however, the inequality simplifies to
@@
$\left|\Pi\!\left(\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\right)-(1-\alpha)\right|\leq\frac{1+\max(\alpha,1-\alpha)}{N}+\sqrt{\frac{1}{2N}\ln\!\left(\frac{2}{\delta}\right)}$

### v2

@@ 3 Conformalized Bayesian Inference: preliminaries and point estimation
Let $\theta,\theta^{\prime}\in\Theta$. The $\mathcal{D}$-KDE score $s(\theta;\boldsymbol{\theta}^{1})$ with hyperparameter $\gamma>0$ is given by
$s(\theta;\boldsymbol{\theta}^{1}):=\frac{1}{S}\sum_{t=1}^{S}\mathcal{K}(\theta,\theta_{t}),\qquad\mathcal{K}(\theta,\theta^{\prime}):=e^{-\gamma\mathcal{D}(\theta,\theta^{\prime})}.$
Intuitively, $s(\cdot;\boldsymbol{\theta}^{1})$ acts as a KDE score, but with the Euclidean norm replaced by the metric $\mathcal{D}$ to handle the possibly complex parameters in $\Theta$ (scholkopf2001learning, shawe2004kernel). Under certain conditions on $(\Theta,\mathcal{D})$, one can have $\mathcal{D}(\theta,\theta^{\prime})=\|\phi(\theta)-\phi(\theta^{\prime})\|$ for some isometric embedding $\phi:\Theta\to\mathbb{R}^{E}$ and $E\in\mathbb{N}$, at least for certain $\theta,\theta^{\prime}$ (schoenberg1935remarks). Thus, the proposed score can be viewed—formally in some cases, loosely in general—as a two-step procedure: first embedding parameters into a latent Euclidean space that preserves the geometry induced by $\mathcal{D}$, and then performing KDE in that space.
While other scoring rules can be devised, $\mathcal{D}$-KDE provides a simple and robust baseline. [^5] A kernel different from $D\mapsto e^{-\gamma D}$ may also be chosen. It relies on a clearly defined metric that encodes a meaningful notion of distance between parameters (a computationally and/or conceptually valuable feature, depending on the context) and reduces the challenge of defining a pseudo-density score on a complex space $\Theta$ to a straightforward, efficiently implementable KDE computation. In particular, evaluating $\mathcal{K}(\theta,\theta_{t})$ is easily parallelized across samples in the calibration and training sets. Optionally, computation can be further sped up by building the score, for each calibration sample, using only a random subset of the training samples.
@@
$\theta_{\star}\in\arg\max_{\theta\in\boldsymbol{\theta}^{2}}s(\theta;\boldsymbol{\theta}^{1}).$
Because $s(\cdot;\boldsymbol{\theta}^{1})$ is constructed to act as a posterior KDE evaluation, $\theta_{\star}$ can be interpreted as a kind of maximum-a-posteriori (MAP) estimator based on this posterior KDE. Of course, this interpretation should be taken with caution, since the posterior may not possess a proper or tractable density (as in random partition models, where the posterior over $\Theta$ is a probability mass function and the true MAP corresponds to the partition with the largest mass), although it provides a simple and efficient strategy to summarize the posterior with a single parameter value.
Another useful perspective is to view $\theta_{\star}$ as a posterior expected loss minimizer restricted to the calibration set, with the only difference that the loss $\mathcal{D}$ is transformed via $D\mapsto\exp\{-\gamma D\}$ to produce a pseudo-density score. In contexts such as random partition models, where efficient algorithms exist for exhaustive searches over $\Theta$ (wade2018clustering, dahl2022search), ideas from that literature can be used to extend the search for the score maximizer beyond the calibration set. Nonetheless, we believe that restricting the maximization to the calibration set remains appealing for several reasons: (i) the scores must in any case be computed (and can be parallelized) for the uncertainty quantification step based on conformal prediction (described later), making this approach more convenient than (possibly sequential) exhaustive search algorithms, especially when the primary goal is uncertainty quantification and point estimation serves only as a simple preliminary summary of the posterior; (ii) the calibration samples are drawn from the posterior and are therefore likely to achieve high scores in the first place, making exhaustive search potentially redundant; and (iii) the search can be flexibly customized—for example, in the context of clustering, one may prefer focusing on partitions with at most three scientifically interpretable clusters, in which case the score maximization can be restricted to the subset of calibration samples satisfying the desired properties. This allows the procedure to combine data-driven evidence with user-specified structural assumptions in a flexible and interpretable way.
@@
The data are then analyzed by fitting a Gaussian mixture model with a Pitman-Yor process prior on the mixing measure (pitman1997two, perman1992size, deblasi2013gibbs), using a concentration parameter of 0.03 and a discount parameter of 0.01. A standard marginal Gibbs sampler is applied via the Python package pyrichlet (selva2025pyrichlet), and after burn-in and thinning, 5,000 training partitions and 1,000 calibration partitions are retained.
@@
Figure 1(b) shows the posterior distribution of $K$, the number of distinct clusters to which the 100 data points are assigned. The histogram suggests that the values 2, 3, and 4 all receive substantial posterior mass, indicating the presence of potential uncertainty in the clustering. Figure 1(c) displays the partition corresponding to the $\mathcal{D}_{VI}$-KDE point estimate (with hyperparameter $\gamma=0.5$). This estimate essentially coincides with the true clustering once the two left-most clusters are collapsed together. [^6] In the rest of the paper, we will refer to the partitions pictured in Figures 1(a) and 1(c) as the “true partition” and the “collapsed partition,” respectively. Notably, this occurs even though partitions with 3 clusters receive more posterior mass than those with 2 clusters (see Figure 1(b)). This behavior may be due to many sampled partitions with 3 clusters being very close to the point estimator (e.g., differing from it by only a few cluster assignments), or because the VI metric inherently favors parsimonious clusterings by weighting larger clusters more heavily; see Equation (1). A clearer picture will emerge with our upcoming discussion of posterior uncertainty quantification beyond point estimation.
4 Credible region
@@
While point estimation provides a useful first summary of the posterior distribution, it necessarily omits the uncertainty encoded in the full posterior. To address this, a key component of our CBI methodology is the construction of a credible posterior subset of $\Theta$ based on the principles of conformal prediction (vovk2005algorithmic, shafer2008tutorial, lei2018distribution, angelopoulos2023conformal). Very briefly, conformal prediction in its basic form allows one to construct a set-valued prediction for the next observation given the first $N$ samples from an exchangeable sequence, with a precisely defined, finite-sample, and distribution-free coverage guarantee. In our context, the crucial step is to interpret Monte Carlo samples as draws from an exchangeable sequence, and to recast the problem of constructing a posterior credible region as that of predicting a region where a new independent posterior sample would fall with high probability. [^7] We note that, although with entirely different goals (namely, calibrating predictions derived from Bayesian predictive distributions), conformal prediction has already been applied in the Bayesian setting, e.g., by fong2021conformal. We now formalize this intuition.
@@
Recall our assumption that the available samples from $\Pi$, and in particular the calibration samples $\boldsymbol{\theta}^{2}=(\theta_{S+1},\ldots,\theta_{T})$ (for convenience, let N:=TSN:=T-S), are iid. See Proposition 2 for a theoretical justification of this assumption when the original sampler features Markov dependence, as in MCMC. Define
$\hat{p}(\theta;\boldsymbol{\theta}^{2}):=\frac{|\{\theta^{\prime}\in\boldsymbol{\theta}^{2}:s(\theta;\boldsymbol{\theta}^{1})\geq s(\theta^{\prime};\boldsymbol{\theta}^{1})\}|+1}{N+1}.$ (2)
@@
$\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2}):=\{\theta\in\Theta:\hat{p}(\theta;\boldsymbol{\theta}^{1})\geq\alpha\}.$ (3)
In words, we declare $\theta\in\Theta$ to belong to the credible region $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$ if its $\mathcal{D}$-KDE score is not too low compared to those of the calibration samples—that is, if it is not too “unlikely” under the posterior relative to samples known to be drawn from it. Note that we emphasize the dependence of the credible region on the calibration set but not on the training set. This is because, from now on, the latter is treated as fixed (or, equivalently, we condition on its realization), since the coverage guarantees to be discussed do not depend on any particular scoring rule. The following proposition states precisely the coverage property ensured by $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$.
@@
The set $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$ satisfies
@@
$\mathbb{P}_{(\boldsymbol{\theta}^{2},\theta)\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\theta\in\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})\right]\,\equiv\,\mathbb{E}_{\boldsymbol{\theta}^{2}\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\Pi\left(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})\right)\right]\,\geq\,1-\alpha.$ (4)
Furthermore, for any $\delta\in(0,1)$, its posterior mass satisfies
@@
$\left|\Pi\left(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})\right)-(1-\alpha)\right|\leq\frac{\max(\alpha,1-\alpha)}{N}+\left|\frac{k_{\alpha,N}}{N}-\hat{\Pi}_{s,N}\big(s_{(k_{\alpha,N})}^{-}\big)\right|+\sqrt{\frac{1}{2N}\ln\left(\frac{2}{\delta}\right)}$ (5)
@@
with $\mathbb{P}_{\boldsymbol{\theta}^{2}\overset{\textnormal{iid}}{\sim}\Pi}$-probability at least $1-\delta$, where $k_{\alpha,N}:=\lceil\alpha(N+1)-1\rceil$, $s_{(k)}$ is the $k$-th order statistic of the calibration scores, and $\hat{\Pi}_{s,N}$ is their empirical cdf. Finally, $\hat{p}(\theta;\boldsymbol{\theta}^{2})$ satisfies
@@
$\forall\lambda\in[0,1],\quad\mathbb{P}_{(\boldsymbol{\theta}^{2},\theta)\,\overset{\textnormal{iid}}{\sim}\,\Pi}\left[\hat{p}(\theta;\boldsymbol{\theta}^{2})\leq\lambda\right]\leq\lambda.$ (6)
@@
Proposition 1 clarifies the rationale behind defining $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$ as a credible region under the posterior $\Pi$ with coverage $(1-\alpha)\times 100\%$. In particular, Equation (4) ensures that the entire procedure, i.e., scoring the calibration samples and checking whether an independent draw from $\Pi$ falls within $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$, succeeds with probability at least $1-\alpha$. Unsurprisingly, given its connection to conformal prediction, this construction has a frequentist flavor: it yields a distribution-free, finite-sample coverage guarantee under $\Pi$ based on samples from it. From a purely subjective Bayesian standpoint, however, this raises no philosophical tension, since the task is to infer properties of the posterior distribution itself, which exists and is fixed, independently of whether the Bayesian model giving rise to it is meant to learn any true, fixed data-generating process. On the other hand, a key advantage of the conformal prediction approach is that the resulting coverage guarantee is not tied to any property of the posterior distribution. In particular, it does not rely on any specific topological or algebraic structure of $\Theta$, which in many cases of interest may be non-standard, and remains valid regardless of the dimensionality of the posterior support, which in many of our motivating applications is extremely large or even infinite.
@@
As Equation (4) shows, the coverage guarantee can also be interpreted as an expected-mass statement: the posterior mass of $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$ exceeds $1-\alpha$ in expectation, where the randomness arises from the calibration set. Thus, $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})$ is a random subset of $\Theta$ whose expected posterior mass is at least $1-\alpha$. Equation (5) further characterizes the behavior of this random set, showing that, with high probability, its posterior mass concentrates around $1-\alpha$ at rate $N^{-1/2}$ in the calibration size $N$. Note that, unless the posterior distribution of the scores is atomless, the second term on the right-hand side of Equation (5) need not vanish as $N\to\infty$ (while it equals $1/N$ when the score distribution is continuous). This reflects the well-known fact that conformal quantile-based sets achieve coverage at least as high as the nominal level $1-\alpha$, but may not approximate it tightly from above if atoms at specific quantiles exist in the target score distribution.
@@
This notion of coverage differs from the traditional Bayesian definition, which concerns a fixed, non-random credible region $\mathcal{C}$ satisfying $\Pi(\mathcal{C})\geq 1-\alpha$. Such a region, if computable and sufficiently small, would of course be preferable, as it would provide deterministic coverage. However, the very complexity of $\Theta$ that motivates our framework typically precludes analytical construction of such sets. Moreover, even conventional credible regions derived from Monte Carlo samples are inherently random, although this is rarely acknowledged explicitly: for instance, even a simple credible interval for a one-dimensional parameter (e.g., based on empirical upper- and lower-quantiles of Monte Carlo samples from the posterior distribution of that parameter) inherits randomness from the finite Monte Carlo sample size. From this point of view, a key advantage of our conformal construction is that it delivers rigorously defined, finite-sample posterior coverage guarantees, both in expectation (Equation (4)) and with high probability (Equation (5)), that hold for any calibration size $N$ and without any assumption at all on the target posterior distribution.
@@
Moreover, Equation (6) shows that $\hat{p}(\theta;\boldsymbol{\theta}^{2})$ can be interpreted as a $p$-value for testing the null hypothesis that $\boldsymbol{\theta}^{2}$ and $\theta$ are drawn iid from $\Pi$, since its null distribution is super-uniform. This enables formal hypothesis testing under the posterior, though of a rather different nature from traditional Bayesian procedures such as tests relying on Bayes factors (jeffreys1998theory, kass1995bayes). The proposed conformal test has a distinctly frequentist flavor, as it treats the posterior $\Pi$ as an unknown distribution estimated from iid Monte Carlo samples, and the conformal $p$-value provides a classical tool to test a nonparametric null hypothesis on the joint distribution of the calibration samples and a hypothetical new draw from $\Pi$. More generally, if one prefers not to attach a strict testing interpretation to $\hat{p}(\theta;\boldsymbol{\theta}^{2})$, this quantity can still be viewed as a normalized measure of the typicality of $\theta$ with respect to $\Pi$: larger values correspond to parameters that rank higher, under the $\mathcal{D}$-KDE score, than most calibration samples. Even under this heuristic interpretation, the conformal analysis based on $\hat{p}(\theta;\boldsymbol{\theta}^{2})$ offers a more nuanced characterization of the uncertainty encoded in $\Pi$ than a simple point estimate alone could provide.
@@
An additional appealing feature of the procedure is that it may be applied conditionally on any subregion of $\Theta$ with positive posterior mass. This is particularly useful when one wishes to perform inference while fixing certain aspects of the complex parameter. For example, as already mentioned, in clustering applications one may wish to assess posterior uncertainty conditional on partitions with at most three interpretable clusters. In this case, a valid credible region can be constructed from the calibration scores associated with partitions having at most three clusters, with coverage guaranteed relative to the posterior conditioned on the nonzero-$\Pi$ event of no more than three clusters.
@@
On a final note on Proposition 1, it is well known (angelopoulos2023conformal) that the coverage guarantee in Equation (4) is indepenent of the choice of scoring rule, as it relies solely on the exchangeability between calibration and new samples. Nevertheless, the practical utility of the resulting credible region depends critically on this choice. Arbitrary scores may yield overly large regions (since even the whole $\Theta$ itself trivially satisfies the guarantee) or exclude high concentration areas if the score misaligns with $\Pi$. Designing an informative score is therefore essential. Our distance-based KDE construction offers a principled solution: it exploits (i) the geometry of $\Theta$ induced by $\mathcal{D}$, which helps capture the structure of the effective posterior support, and (ii) the inherent concentration of training samples in high-probability regions. In our illustrations, we verify empirically that this choice yields credible regions of reasonable size and consistent posterior alignment in practice.
@@
The concentration inequality in Equation (5) is of independent interest, and its original proof, based on the Dvoretzky–Kiefer–Wolfowitz inequality (dvoretzky1956asymptotic, massart1990tight), can be found in the Supplementary Material. It is in the spirit of existing results (vovk2012conditional, sarkar2023post, lei2018distribution) that study the coverage of the conformal set as a random variable conditional on the calibration scores, and establish high-probability bounds for the event of that coverage exceeding $1-\alpha$. Our inequality goes a step further (under the assumption of iid scores), as it specifies when and at what rate this coverage converges exactly to $1-\alpha$, rather than merely exceeding it. As already noted, the bound includes a term that may not vanish asymptotically in the presence of atoms in the score distribution; when $\Pi$ is continuous, however, the inequality simplifies to
@@
$\left|\Pi\!\left(\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})\right)-(1-\alpha)\right|\leq\frac{1+\max(\alpha,1-\alpha)}{N}+\sqrt{\frac{1}{2N}\ln\!\left(\frac{2}{\delta}\right)}$